### v2

# Curiosity-Driven Co-Development of Action and Language in Robots Through Self-Exploration

Abstract
Human infants acquire language and action co-developmentally, achieving remarkable generalization capabilities from only a minimal number of learning examples. In contrast, recent large language models require exposure to billions of training tokens to achieve such generalization. What mechanisms underlie such efficient developmental learning in humans? This study addresses this question through simulation experiments in which robots learn to perform various actions corresponding to imperative sentences (e.g., push red cube) via trials of self-guided exploration. Our approach integrates the active inference framework with reinforcement learning, enabling curiosity-driven developmental learning. The simulations yielded several nontrivial findings: i) Curiosity-driven exploration combined with motor noise substantially outperforms learning without curiosity. ii) Simpler, prerequisite-like actions emerge earlier in development, while more complex actions involving these prerequisites develop later. iii) Rote pairing of sentences and actions occurs before the emergence of compositional generalization. iv) Generalization is drastically improved as the number of compositional elements increases. These results shed light into possible mechanisms underlying efficient co-developmental learning in infants and provide computational parallels to findings in developmental psychology.
Introduction
A central question in both cognitive science and artificial intelligence is how humans and artificial systems can acquire competencies for language and motor command in a co-developmental manner, despite having access to only limited learning experiences. This question is exemplified in human infants, who achieve remarkable generalization with sparse input. This is a stark contrast to large-scale models which rely on massive training corpora, to reach similar capabilities. This raises the issue of what mechanisms enable such efficient developmental learning.
From the perspective of developmental psychology, infants acquire language through rich interaction with their embodied environments. Tomasello’s “verb-island” hypothesis argues that children initially learn verbs in specific, isolated contexts before generalizing across broader linguistic structures (?). He also emphasized the importance of embodiment in language acquisition, suggesting that grounding linguistic symbols in sensorimotor experiences is fundamental to language learning (?). This view aligns with other studies in developmental psychology highlighting the role of compositionality and generalization in language acquisition (?, ?, ?).
In linguistic terms, compositionality refers to the ability to construct novel configurations by systematically combining elements such as verbs, adjectives, and nouns. Generalization enables infants to apply learned components flexibly, allowing for the production and interpretation of utterances that have not been directly encountered previously. Although the number of possible compositions grows multiplicatively with the vocabulary size (i.e., number of verbs $\times$ number of adjectives $\times$ number of nouns), infants achieve generalization after experiencing only a small subset of learning examples. This suggests that the effective sample complexity is closer to the sum of elements rather than their product. This phenomenon is closely related to the “poverty of the stimulus” problem articulated by Chomsky (?), which asks how learners generalize so effectively given severely sparse input.
To investigate these mechanisms, one promising approach is to reconstruct developmental learning processes in machines and robots. The field of developmental robotics has long pursued this line of research, aiming to replicate human-like learning trajectories in embodied systems (?, ?, ?, ?). However, relatively few studies have focused on the co-development of language and motor control under conditions of stimulus poverty. Existing work has primarily examined associative mappings between linguistic input and motor commands in one-shot or supervised batch learning schemes (?, ?, ?, ?). These approaches neglect the self-directed, developmental context of infant learning.
In this study, we propose a self-exploratory learning framework of robots in which reinforcement learning is incorporated with the active inference framework (?, ?, ?), enabling curiosity-driven exploration. Our approach to integrate reinforcement learning with active inference was originally inspired by the work of Kawahara et al. (?). In our model, originally introduced in (?), motor commands are reinforced by two intrinsic rewards: curiosity (seeking unpredictable sensory consequences) and motor entropy (seeking random movements). Motor commands are also reinforced by extrinsic rewards for successfully achieving goal tasks specified by given imperative sentences. Importantly, our previous experiments in maze navigation demonstrated that the combination of curiosity and entropy is crucial for enhancing self-exploration, as agents achieved significantly improved exploratory behaviors under this dual-intrinsic reward scheme. Our approach aligns with broader research on self-exploration in machine learning, in which agents are intrinsically rewarded for taking motor commands that increase unpredictability or information gain (?, ?).
A simulated mobile robot equipped with a manipulator arm, vision sensor, and distributed tactile sensors learns to generate motor movements in response to imperative sentences presented during each trial. These sentences are systematically composed of verbs, nouns, and adjectives, enabling evaluation of generalization performance under different levels of compositional complexity.
The model architecture employed in this study is based on our previous work (?) with key modifications to accommodate multi-modal sensorimotor integration. (See details in the Materials and Methods section.) Figure 1 presents the model architecture, which is composed of two main components: a forward model and an actor part.
Figure 1: The proposed model architecture. The model consists of predictor, actor, and critic.
The forward model learns to predict the next sensation $o_{t+1}$ based on the current sensation $o_{t}$ and the executed motor command $a_{t}$. The sensation includes pixel-based vision, tactile sensation, arm joint proprioception, and voice for the sentences. To address the hidden state problem and probabilistic nature of the environment, the prediction is performed contextually and stochastically using the random latent variables $z_{t}$ and the deterministic $h^{q}_{t}$. The random latent variables were allocated separately for each sensory modality, while the deterministic variables were shared. The actor module generates the next motor command $a_{t+1}$ based on the deterministic latent variable $h^{q}_{t}$, which integrates current sensation $o_{t}$.
The overall flow is: with an imperative sentence given by the tutor, the robot attempts to achieve the specified goal by generating a sequence of motor commands. Meanwhile, the forward model predicts the next sensation by inferring the posterior probability distribution $q(z_{t}|o_{t},h_{t-1})$ of the random latent variable $z$. This inference is conducted by minimizing the evidence free energy $F$ (Eq. 1). This consists of the complexity term represented by Kullback–Leibler divergence (KLD) between the estimated posterior and the prior, and the accuracy term as shown in the free energy principle (FEP). (See more details on the evidence free energy and expected free energy in the “Free Energy Principle, Active Inference, and Kawahara Model” section of the Supplementary Materials.)
$F_{\psi,t}=\underbrace{D_{KL}[q(z_{t}|o_{t},h_{t-1})||p(z_{t}|h_{t-1})]}_{\text{Complexity}}-\underbrace{\mathbb{E}_{q(z_{t})}[\log p(o_{t+1}|h_{t})]}_{\text{Accuracy}}.$ (1)
The forward model is trained incrementally by optimizing the learning parameters $\psi$ in the direction of minimizing the evidence free energy. The actor generates motor command sequences in the direction of minimizing the expected free energy $G$ (Eq. 2). This consists of the complexity term, extrinsic reward term, and the entropy term through reinforcement learning of motor commands.
$G(a_{t})$ $=-\underbrace{D_{KL}[q(z_{t}|o_{t},h_{t-1})||p(z_{t}|h_{t-1})]}_{\text{complexity}}-\underbrace{r(s_{t},a_{t})}_{\text{Extrinsic Reward}}-\underbrace{\mathcal{H}(\pi_{\phi}(a_{t}|h_{t-1}))}_{\text{Entropy}}$ (2)
It is interesting to note that minimizing evidence free energy $F$ minimizes the complexity term while minimizing expected free energy $G$ maximize the same complexity term. This means that motor commands are generated in the direction of maximizing the information gain represented by KLD between the estimated posterior and the prior after the motor command execution. This generates curiosity-driven exploration wherein the agent seeks out previously unencountered sensorimotor experiences. On the other hand, the learning parameters in the forward model is updated in the direction of minimizing the same complexity term, representing the latent conflict that is generated by novel experiences encountered during curiosity-driven exploration. Therefore, both processes of self-exploration and the forward model learning are racing each other.
This study tested the following hypotheses through simulation experiments: H1: Curiosity combined with motor entropy enhances the performance of developmental learning. H2: Primitive actions are acquired earlier, followed by more complex, prerequisite-dependent actions. H3: In the early phase, actions are generated only for exactly learned sentences, but in later phases, the system generalizes to novel, unlearned compositions. H4: Generalization performance improves as the scale of compositionality in the task increases.
Results
Task Description
We created a robot like a truck crane in a physics simulator along with a set of objects with 5 different shapes each of which can be with 6 different colors (see Figure 2).
Figure 2: The simulated robot and a set of objects to act on. Top: the robot has two wheels and an arm with two joints. The design is similar to a truck crane. Bottom, left to right: a red pillar, a green pole, a blue dumbbell, a cyan cone, and a magenta hourglass. The color yellow is not pictured here.
The robot can maneuver by controlling velocity of left and right wheels independently, and also can move its arm by controlling rotation velocity of the yaw and pitch joint angles for acting on the objects. A camera with 16 x 16 pixels was fixed to the body for visual sensation. 16 touch sensors were distributed in the body and the arm, respectively, and rotation angles for the yaw and pitch were sensed as proprioception.
For each trial episode, a task goal was given in terms of an imperative sentence composed with verb, adjective, and object noun. Possible words used for them are shown in Table 1.
English Words
Verb Adjective Object noun
watch red pillar
be near green pole
touch the top blue dumbbell
push forward cyan cone
push left magenta hourglass
push right yellow
Table 1: The English words used for imperative sentences specifying goals.
In the beginning of each episode, two objects were located at random positions in the arena wherein one object was the one specified in the imperative sentence and the other was the one with randomly selected color and shape combination among possible ones.
At each step, the robot receives visual sensation, proprioception for the arm, tactile sensation , and two types of voices: the command voice and the tutor-feedback voice. The command voice takes the format of the imperative sentence described previously, and it comes every step continuously from the beginning. On the other hand, the feedback voice arrives whenever the robot achieves one of possible goals even if the achieved goal is not the one told by the command voice, and it informs which goal has been achieved actually in the same format with the command voice. This potentially enhances the forward model to learn about own motor motions. Finally, when the goal specified by the command voice is achieved, a reward is provided. Each trial episode ran for 30 steps, or terminated when the specified goal is achieved.
Effects of curiosity: Experiment 1
The experiment examined effects of different levels of curiosity to the developmental learning processes using the basic setup. In the basic setup, full compositions of words (Table ) were used to generate the imperative sentences. However, the training was conducted using only 60 imperative sentences (33 percentage) out of 180 possible sentences. 120 untrained sentences were used for generalization test. For ten agents with different random seeds, the whole developmental learning process was iterated for 60000 epochs. The generalization test with unlearned imperative sentences was conducted for every 50 epochs.
The experiment was conducted by changing the levels of curiosity. Since the random latent variables are computed separately for each sensory modality, the complexity or curiosity can be computed for each sensory modality. Three levels of curiosity were considered in computing expected free energy $G$ wherein no curiosity: the curiosity terms for all sensory modalities are not included, sensory-motor curiosity: the curiosity terms only for vision, tactile sensation, and proprioception are included, all curiosity: the curiosity terms for all sensory modalities including feedback voice are included.
Figure 3 shows the development of the generalization test performances in terms of success rate for goals specified by unlearned imperative sentences which are plotted for different action categories with different levels of curiosity. Shades areas represent 99% confidence intervals.
Figure 3: Rolling success-rates for unlearned goals for agents with different levels of curiosity.
The plots shows that the performance was improved significantly as the curiosity level was increased. Especially, the case of all actions with the all curiosity level shows the average success rate for unlearned goals reached a quite high value of 90 percentage even though the learning was conducted only for 33 percentage of all possible compositions. (See a video: https://www.youtube.com/watch?v=ipjXg53f-zk) showing examples of the robot behaviors with the all curiosity level in the intermediate phase and the final phase in the course of development. It can be also seen that some action categories developed faster than other action categories. Especially in the all curiosity case, “watch” developed the fastest, “be near” did as the second, “push forward” as the third, and “push left,” “push right” and “touch the top” developed much later. This implies that simpler prerequisite-type actions develop earlier, and more complex actions requiring those prerequisite actions develop later. Actually, an action of watching an object should be prerequisite for all other object-targeted actions including an action of moving near by an object, which should be prerequisite again for actions of directly manipulating an object like pushing left/right an object or touching the top of it. Our observation accords with this.
Next, Figure 4 (A) shows the success rate comparison between learned and unlearned goals under the all curiosity condition for each action category. These plots show that the test performance for learned goals developed significantly faster than the case for the unlearned goals. This indicates that actions are generated only for exactly learned goal compositions in the early phase, but the system generalizes to novel, unlearned composition ones in the later phase.
Effects of scale in compositions: Experiment 2
Next experiment examines the effects of scales of compositionality in learned examples to the generalization performance. For this purpose, experiments were conducted using reduced number of words for generating imperative sentences. While the previous basic setup used sentences composed of 6 verbs, 6 adjectives, and 5 object nouns as the full scale case, the middle scale case was prepared with 5 verbs, 5 adjectives, and 4 object nouns, and the small scale case with 4 verbs, 4 adjectives, and 3 object nouns. The exact words used for each setup are listed in Table 4. For all scaling cases, again only one third was used for learning examples while remained two third was used for generalization test. Other experimental conditions were also set as the same as the Experiment 1.
The experimental results are shown in Figure 4. Shades areas represent 99% confidence intervals.
Figure 4: Rolling success-rates for agents with different numbers of tasks, colors, and shapes.
It can be seen that although the learned goal test cases show equally high performance for all scales of compositionality, the generalization test for unlearned goal case shows significantly lower performance as the compositionality scale decreases. This indicates that the generalization performance severely depends on the scale of compositionality in learning examples.
Further analysis on the experiment results
Some analysis were conducted for the purpose of gaining comprehension of the internal representation developed. We applied Principle Component Analysis (PCA) to the estimated posterior latent states corresponding to the command voice input, incorporating the tasks and colors stated in each command. Figure 5 presents evidence that robots with all curiosity developed a a compositional and generalizable understanding of these goals. At the midpoint of training, the latent representations begin to show consistent grouping of tasks and colors. For example, the tasks “watch,” “be near,” and “push forward” are tightly grouped, suggesting that these tasks are interpreted as similar. In contrast, the task “push right” appears heavily separated from other tasks, and the tasks “touch the top” and “push left” also appear as distinct categories. After training, these clusters of tasks are more compact and separated. Within these clusters, there is loose sub-structuring by color: green and yellow tend to be on the left side of the cluster, while blue and magenta are on the right side of the cluster. Therefore, it can be said that each cluster represents a distinct linguistic concept while exhibiting relationship with others since the learning of visuo-proprioceptive-motor also contributes to this structuring. In contrast, Figure 6 shows the same type of PCA results for a robot with no curiosity. “Be near,” “push forward,” and “push right” are heavily entangled with each other; coincidentally, in Figure 3, these are the three tasks which these robots perform worst. Thus supports the idea that curiosity aids task disentanglement and compositional learning.
Figure 5: Compositional understanding in robots with all curiosity. PCA of latent variables corresponding to the command voice. Left: halfway through training. Right: after training. Clusters with substructures emerged early and became more refined over time.
Figure 6: Compositional understanding in robots with no curiosity. PCA applied to latent representations of command voice inputs after training. Tasks with poor performance remain entangled.
In the current model, the robot’s knowledge in the environment should develop richer along the course of exploratory learning. For confirmation of this idea, we examined the capability of the robots in generating mental plans for achieving goals without accessing the sensory inputs except the initial step for an episode trial as compared between the half-trained case and the fully trained case. A robot’s mental planning can be visualized by allowing it to receive real sensory observation only at the initial step, after which the robot must rely entirely on its own internal predictions. In this setting, the robot views its predicted sensory observations as if they are true inputs. This process may be likened to a dreamlike state or hallucinatory simulation, in which the robot mentally simulates future events based on its internal model of the world. Figure 7 (a) illustrates a simulation example for the fully trained case. The robot was commanded to touch the top of the yellow pillar. In that figure, the first row shows the ground truth environment from a view behind the robot’s shoulder. The second row shows what the robot would truly observe if it were not in this planning setting. The third row shows the robot’s look ahead predictions for visual observations, which it will interpret as if they are real. It can be seen that even with only the first step sensory observation, the robot could generate mostly accurate future look-ahead prediction for sensation as well as motor command. These predictions are sufficiently accurate for the robot to maintain an internal conceptualization of the environment and complete its command in the case of the end of the developmental learning. Figure 7 (b) illustrates the same robot after only half of its training in the same scenario. In this case, the robot’s predictions are inaccurate, causing it to wander and view an object which does not actually exist.
Figure 7: A mentally planning robot commanded to touch the top of the yellow pillar. The first row displays the ground truth from a view over the robot’s right shoulder. The second row displays the vision which the robot would have experienced if it were not only mentally planning; the robot observes only the first genuine sensory observation. Subsequent inputs are the robot’s own predictions, which it treats as real, displayed in the third row. (A) Despite the limitation, the robot successfully completes the commanded action and maintains a consistent internal model of its environment. (B) The same robot in the same scenario, but after only half of its training, is unable to make correct predictions and therefore cannot complete the command action.
Discussion
This study investigated how robots can co-develop action and language through self-exploration by integrating active inference with reinforcement learning. The experiments were designed to test four specific hypotheses, and the results provide clear support for each.
H1: Curiosity combined with motor entropy enhances developmental learning. The first hypothesis was confirmed. Robots equipped with both curiosity-driven exploration and motor entropy consistently outperformed those without, achieving higher success rates in both learned and unlearned tasks. This advantage was particularly pronounced when curiosity extended across all sensory modalities, including vision, touch, proprioception, and voice feedback. These findings suggest that the synergy of curiosity (seeking novel, unpredictable outcomes) and entropy (encouraging stochastic exploration) plays a crucial role in accelerating the acquisition of language-action mappings. This result is consistent with our previous study (?), which showed that combining curiosity and entropy significantly enhanced self-exploration in a maze navigation task. More broadly, this interpretation aligns with the active inference framework, in which agents minimize expected free energy by jointly reducing uncertainty and maximizing information gain (?, ?).
H2: Generalization follows rote learning. The results further supported the second hypothesis. In the early phases, robots were able to perform actions only for sentences encountered during training. Over time, however, they generalized to novel sentences composed of familiar words. This developmental trajectory—rote learning preceding compositional generalization—closely parallels patterns observed in human infants. Tomasello’s “verb-island” hypothesis, for example, emphasizes that children initially acquire verbs in isolated contexts before generalizing across broader structures (?). Such progression resonates with our findings, where early rigid associations gave way to flexible recombination of linguistic and motor elements.
H3: Primitive actions precede complex actions. The third hypothesis was also validated. Simpler, prerequisite-like actions such as “watch” or “be near” emerged earlier, while more complex manipulative actions like “push left” or “touch the top” developed later. This ordering mirrors hierarchical dependencies in action acquisition observed in infant development (?), reflecting how foundational skills scaffold the emergence of more sophisticated behaviors.
H4: Generalization is enhanced by compositional scale. Finally, the experiments confirmed that larger vocabularies of verbs, adjectives, and nouns led to greater generalization. Robots trained with richer compositional repertoires achieved higher success rates on unlearned tasks, whereas smaller vocabularies constrained generalization severely. Our previous study (?) on supervised training of language and action for an arm robot also showed that compositional generalization improved as the size of verb–noun combinations in training increased. However, that work was limited in scale, examining only cases from $3\times 3$ to $5\times 8$ verb–noun combinations, where success rates for unlearned goals improved only from 57% to 71% under the condition of 80% training. By contrast, the present study examined much broader scaling, ranging from 48 to 180 possible compositions, while using only 33% of them for training. Under these conditions, generalization performance improved dramatically from near 25% to 90%. This contrast highlights that scale plays a critical role in enhancing generalization, and that curiosity-driven developmental learning provides a more powerful mechanism than supervised schemes under conditions of limited input. The finding also connects to the classical “poverty of the stimulus” problem raised by Chomsky (?), as it shows how compositionality enables powerful generalization from sparse training data. Once, we hypothesize that necessary training size could be proportional to summation of number of words appeared for each dimension instead of multiplication of it for all dimensions if compositionality size increases (?). This hypothesis becomes more plausible by the results in the current study which, however, should be confirmed in much more scaled experiments in the future.
Taken together, these findings demonstrate that curiosity-driven exploration, motor entropy, hierarchical acquisition of actions, and scalable compositional exposure jointly support efficient co-developmental learning of language and action. The parallels with infant development—rote-to-generalization progression, prerequisite learning, and the role of vocabulary scale—suggest that the mechanisms implemented here capture essential aspects of developmental psychology. More broadly, these results strengthen the view that reconstructing developmental processes in robots can offer insights into the “poverty of the stimulus” problem, showing how powerful generalization can arise from limited input when guided by intrinsic motivation, structured experience, and the principles of predictive coding and active inference (?, ?, ?, ?).
Future Directions
The current study is still limited in many aspects, and several possible extensions can be envisioned. One crucial limitation is that our experiments examined only a one-directional communication pathway from tutors to robots, relying on the command and feedback voices to guide the development of language–action mappings. In contrast, natural human development is characterized by interactive and bidirectional communication, where infants not only receive instructions but also actively solicit guidance, clarification, and scaffolding from caregivers.
Future studies should extend the current framework to include interactive communication between tutors and robots. For example, when a robot cannot successfully execute a command, it could initiate a communicative act such as “Tell me how to do it” or “Ask me an easier one.” Such exchanges would allow tutors to adapt their teaching strategy dynamically, modulating the complexity of instructions or providing additional cues. This adaptive interaction resonates with Vygotskian ideas of scaffolding and the “zone of proximal development,” where caregivers adjust support according to the learner’s current abilities (?). It also aligns with research in developmental psychology emphasizing the role of joint attention, imitation, and social feedback in language learning (?, ?).
Incorporating interactive dialogue would thus move the current model closer to capturing the social nature of language acquisition in human infants. By embedding mechanisms for robots to both seek help and influence the tutoring process, future work could shed light on how social scaffolding and communicative feedback accelerate the co-development of language and action in natural developmental contexts.
Another promising future direction concerns the development of robot–robot communication through the evolution of language. Previous research has explored this possibility from different perspectives: Steels introduced the framework of “language games” to study the emergence of shared vocabularies among agents (?), Miikkulainen and colleagues investigated the evolution of artificial language through evolutionary reinforcement learning (?), and Taniguchi proposed the emergence of symbols using a collective predictive coding approach (?). While these studies have nicely demonstrated the possibility of emergent communication, they still remain limited in that they mainly achieved the emergence of object labeling or naming, whereas the evolution of action-related language, such as verbs, has been much less explored.
In this context, the current study based on active inference could be extended to address the evolution of dynamic linguistic structures, including verbs. Since our model implements active inference within a variational recurrent neural network, it is naturally suited for capturing temporal and dynamic aspects of action and language. A future extension of this work toward multi-robot interaction under the framework of “collective active inference” may thus provide novel insights into the evolution of embodied language, moving beyond static object labeling toward dynamic and action-oriented communication.
Materials and Methods
In this section, we present the model architecture employed in this study. The current model, as well as our earlier work (?), extends a study by Kawahara et al. (?). That study demonstrated that curiosity-driven reinforcement learning can be achieved by incorporating the framework of active inference (AIF) (?, ?), in which motor behavior is reinforced in the direction which minimizes expected free energy. More details are shown in the “Free energy principle, Active Inference, and Kawahara Model” section of the Supplementary Materials, along with a brief introduction of the free energy principle (FEP) and AIF.
The Employed Model
The current model, as well as our previous one (?), extends the approach proposed by Kawahara et al. (?) by implementing both the forward model and actor-critic using a variational recurrent neural network (VRNN) (?) in order to deal with temporal complexity and stochasticity inherent in robot–environment interactions.
The expected free energy $G$ can be computed as:
$G_{t}$ $=-\underbrace{\eta D_{KL}[q(z_{t}|o_{t},h_{t-1})||p(z_{t}|h_{t-1})]}_{\text{Curiosity/complexity}}-\underbrace{r(s_{t},a_{t})}_{\text{Extrinsic Reward}}-\underbrace{\alpha\mathcal{H}(\pi_{\phi}(a_{t}|h_{t-1}))}_{\text{Entropy}}$ (3)
This equation is derived by replacing $w$, the probabilistic model learning parameter used in Eq. Free energy principle, Active Inference, and Kawahara Model, with $z$, the probabilistic model state. The weighting coefficients $\eta$ and $\alpha$ are introduced to scale the contributions of the curiosity and entropy terms, respectively. The complexity term is computed as Kullback–Leibler divergence (KLD) between the estimated posterior distribution and the prior distribution over the latent variables at each time step. Both distributions are modeled as Gaussian distribution with time-dependent means and standard deviations. The estimated posterior is conditioned on the current sensory observation and the previous hidden state, while the prior is conditioned only on the previous hidden state. The resulting KLD thus reflects the information gain from that sensory observation, which is driven by the motor command executed at the previous time step. Therefore, exploration of more novel situations (i.e., curiosity-driven exploration) tends to result with higher information gain through larger complexity. The entropy in the third term of Eq. 3 reflects the expected uncertainty of the policy, and is computed as the negative expected log-probability of generating a motor command $a_{t}$ conditioned on the hidden state $h_{t-1}$.
By adopting an analogous approach to the Kawahara model, the policy for generating a motor command $a_{t}$ is trained to minimize the expected free energy $G_{t}$ (Eq. 3) through RL using the the Soft Actor Critic (SAC) algorithm (?). Accordingly, the $Q_{t}$ value is updated as:
$Q_{t}$ $=r_{t}+\eta D_{KL}[q(z_{t}|o_{t},h_{t-1})||p(z_{t}|h_{t-1})]+\alpha\mathcal{H}(\pi_{\phi}(a_{t+1}|h_{t}))$
$+$ $\gamma(1-done_{t})\mathbb{E}_{o_{t+1}\sim D,a_{t+1}\sim\pi_{\phi}}[Q_{\bar{\theta}}(o_{t+1},a_{t+1})].$ (4)
The first term $r_{t}$ represents the extrinsic reward. The second term $D_{KL}[q(z_{t}|o_{t},h_{t-1})||p(z_{t}|h_{t-1})]$ is the intrinsic reward for curiosity, scaled by a positive coefficient $\eta$. The third term $\mathcal{H}(\pi_{\phi}(a_{t+1}|h_{t}))$ is the intrinsic reward for motor entropy, scaled by a positive coefficient $\alpha$. The fourth term is the bootstrapped estimate of the next step’s value, $\widehat{Q_{t+1}}$, which is weighted by a discount rate parameter $\gamma\in[0,1]$. The variable $done_{t}$ is zero for all steps except the episode’s final step, where it is set to one. This restrains the definition of $Q_{t}$ to steps within the episode. The critic $Q_{\theta}(o_{t+1},a_{t+1})$ is trained to generate $\widehat{Q_{t}}$, approximation of $Q_{t}$. The target critic $Q_{\bar{\theta}}(o_{t+1},a_{t+1})$ is maintained for stability in the critic’s training. Initially identical to the critic, the target critic is updated via Polyak averaging such that $\bar{\theta}\leftarrow\tau\theta+(1-\tau)\bar{\theta}$ with $\tau\in[0,1]$. The actor $\pi_{\phi}(o_{t})$ is trained to generate motor commands $a_{t}$ which maximize the critic’s predictions of value. To mitigate positive bias, it is common to train multiple separate critics (each with its own target critic) (?). The actor is trained using the minimum predicted value across critics. Our model employs two separate critics.
The forward model is trained dynamically over the course of exploratory learning by optimizing the model parameters $\psi$ to minimize the evidence free energy $F_{\psi}$ (Eq. 1) after each trial episode. The exact implementation of this process is described in the following subsection, Details of the Model Architecture.
Details of the Model Architecture
This subsection explains further details about the model architecture employed in this current study. As noted earlier, the present architecture extends our previous model (?), which is described in the “Free energy principle, Active Inference, and Kawahara Model” section of the Supplementary Materials. The primary extension involves the use of separate random latent variables, encoders, and decoders for each sensory modality. This design allows the model to process multiple types of sensation independently, including vision, tactile input, proprioception, command voice, and feedback voice. In addition, our model uses an encoder for the 4-dimensional motor command, which includes motor velocities for two the robot’s wheels and two joint angles in its arm. The full architecture of the proposed model is shown in Figure 8.
Figure 8: The details of the proposed model architecture.
Computation in this architecture proceeds as follows:
  1. 1.
    The 4-dimensional motor command from the previous time step is fed into the motor command encoder, producing an encoded motor command vector.
  2. 2.
    The prior distribution for the current time step is computed using the encoded motor command vector and the previous hidden state.
  3. 3.
    The sensory observation for each modality is fed through its corresponding encoder, computing its modality-specific encoded vector.
  4. 4.
    The estimated posterior distribution for each modality is estimated using its sensory encoded vector, encoded motor command vector, and the previous hidden state.
  5. 5.
    All posterior vectors from the current time step are concatenated across all modalities, then sampled and combined with the previous hidden state to compute the current hidden state.
  6. 6.
    The motor command for the current time step is generated from the current hidden state using the actor (policy network).
  7. 7.
    The model predicts the next sensory observation for each modality using the current hidden state and the current motor command, passed through the corresponding sensory decoders.
  8. 8.
    The $Q_{t}$ value is updated according to Eq. The Employed Model.
  9. 9.
    If the episode terminates at this step, the episode’s data is saved in a recurrent replay buffer. A batch of information is sampled from the buffer to train the forward model, actor, and critic. See the Supplementary Materials for more details.
Details of the encoders and decoders of each sensory modality (e.g., vision, tactile sensation, etcetera), as well as the motor command encoder, are described in the “Implementation details” section of the Supplementary Materials.
Robot Actions
The robot and the objects were simulated in PyBullet, the python physics simulator. Each wheel’s velocity was bounded within the range of [10,10][-10,10] meters per second. For scale, the robot’s body is a cube measuring 2 meters along each dimension (length, width, and height). The robot’s arm features two joints: yaw, which rotates left and right within a range of $30^{\degree},30^{\degree}]$, and pitch, which rotates forward and upward within a range of $0^{\degree},90^{\degree}].$ For smooth movement, the robot’s wheel and arm velocities were implemented with linear interpolation from the current to the target velocities.
We defined success criteria for each action category, which determined whether or not the robot earned an extrinsic reward. The distance between the robot and an object was measured from the object’s center to the center of the robot’s body. The robot was considered to be “facing the object” when the angular deviation between the robot’s forward direction the line connection it to the object was less than 15 degrees.
Watch: The robot faces the object between 6 and 10 meters of distance. This must be maintained for 6 steps in a row.
Be Near: The robot faces the object with distance of less than 6 meters, without touching the object. This must be maintained for 5 steps in a row.
Touch the Top: The robot’s hand contacts with the object while the center of the hand is at least 3.75 meters above the floor. This must be maintained for 3 steps in a row.
Push Forward: The robot pushes the object farther than .1 meters, with respect to the robot’s facing direction. This must be maintained for 3 steps in a row.
Push Left: The robot pushes the object to the robot’s left farther than .2 meters, while the robot’s wheels have velocities below 5 meters per second (requiring use of the arm). This must be maintained for 3 steps in a row.
Push Right: Same as Push Left, but in the opposite direction.
There are some constraints in rewarding for actions which are described in the “Constraints in Performing Actions” subsection in the Supplementary Materials.
Acknowledgments
Thank you to the Okinawa Institue of Science and Technology (OIST) and colleagues in OIST’s Cognitive Neurorobotics Unit and Neural Computation Unit for supporting this work.
Funding:
TT was funded by OIST graduate school. J.T. was partially funded by the Japan Society for the Promotion of Science (JSPS) KAKENHI, Transformative Research Area (A): unified theory of prediction and action [24H02175],
Author contributions
T.T. and J.T. designed the model and simulation. T.T., K.D., and J.T. designed the experiments. T.T. performed all simulation experiments and data analisus. T.T. wrote the paper. K.D. and J.T. edited the paper. J.T. supervised the study.
Competing interests:
There are no competing interests to declare.
Data and materials availability:
All files related to this project are available from a repository at the Okinawa Institute of Science and Technology.
The code used in the experiments is accessible via https://github.com/oist-cnru/curiosity_driven_co_development.
The authors are grateful for the help and support provided by colleagues in OIST’s Cognitive Neurorobotics Research Unit and Neural Computation Unit.
Supplementary Materials for
Curiosity-Driven Co-Development of Action and Language in Robots Through Self-Exploration
Theodore Jerome Tinker, Kenji Doya, Jun Tani
Cognitive Neurorobotics Unit, OIST, Onna 904-0495, Japan.
Neural Computation Unit, OIST, Onna 904-0495, Japan.
* To whom correspondence should be addressed; E-mail: jun.tani@oist.jp.
This PDF file includes:
Supplementary Text
Figures S1 to S3
Tables S1 to S8
Other Supplementary Materials for this manuscript:
Supplementary Text
For future reference, table S1 includes definitions for all relevant variables.
Variable Definition
$o_{t}$ Observation at time $t$
$o_{t,i}$ $i^{th}$ part of observation $o_{t}$
$o_{t,v}$ Our agent’s $o_{t,0}$, vision
$o_{t,ta}$ $o_{t,1}$, touch
$o_{t,p}$ $o_{t,2}$, proprioception
$o_{t,cw}$ $o_{t,3}$, command voice
$o_{t,fw}$ $o_{t,4}$, feedback voice
$a_{t}$ Action
$r_{t}$ Extrinsic reward
$done_{t}$ Final step of episode
$mask_{t}$ Steps inside episode
$R$ Recurrent replay buffer
$\pi$ Actor
$\phi$ Actor’s parameters
$Q$ Critic
$\theta$ Critic’s parameter
$\bar{\theta}$ Target critic’s parameter
$\tau$ Critic’s soft update coefficient
$f$ Forward model
$\psi$ Forward model parameters
$\gamma$ Discount for future rewards
$\alpha$ Importance of entropy
$\eta$ Importance of curiosity
$\eta_{i}$ $\eta$ for $i^{th}$ part of observation
$p(z_{t})$, $q(z_{t})$ Prior, estimated posterior
$\mu$, $\sigma$ Mean, standard deviation
$h_{t}$ RNN hidden state
$z_{t}$ Sample from posterior
Variable Definition
$enc_{i}$ Encoder for $o_{t,i}$
$\psi^{enc}_{i}$ $f$ parameters for $enc_{i}$
$dec_{i}$ Decoder for $o_{t,i}$
$\psi^{dec}_{i}$ $f$ parameters for $dec_{i}$
$MLP^{prior}_{i}$ Multilayer for prior for $o_{t,i}$
$MLP^{post}_{i}$ Multilayer for estimated posterior for $o_{t,i}$
Key5 Value10
Key5 Value10
Key5 Value10
Key5 Value10
Key5 Value10
Key5 Value10
Key5 Value10
Key5 Value10
Key5 Value10
Key5 Value10
Key5 Value10
Key5 Value10
Key5 Value10
Key5 Value10
Key5 Value10
Key5 Value10
Key5 Value10
Key5 Value10
Key5 Value10
Key5 Value10
Table S1: Definitions of variables.
Free energy principle, Active Inference, and Kawahara Model
We begin by describing predictive coding and active inference (AIF), which are grounded in the free energy principle (FEP) (?). The FEP posits that biological and artificial agents maintain their existence by minimizing variational free energy, which is an upper bound on sensory surprise. In perception, this process is often instantiated as predictive coding (?, ?, ?, ?), wherein internal models reconstruct sensory inputs by updating beliefs or latent variables by minimizing the reconstruction errors. More formally, this is minimizing evidence free energy defined for past observations. In motor command generation, the FEP framework extends to AIF (?, ?), where agents minimize the future prediction error (quantified as expected free energy) by optimizing the latent variables and motor commands in the future. These two processes are tightly coupled and must be considered jointly in embodied cognition systems.
We next introduce the work of Kawahara et al.  (?), who proposed a novel reinforcement learning (RL) scheme that integrates (AIF).
In the Bayesian framework, the true posterior probability distribution $p(z_{t}|o_{t})$ over latent variables $z_{t}$, conditioned on sensory observations $o_{t}$, is given by Bayes’ rule:
$p(z_{t}|o_{t})=\frac{p(o_{t}|z_{t})p(z_{t})}{\int{p(o_{t},z_{t})dz}}$
Here, $p(z_{t})$ denotes the prior. The denominator, called the evidence, is usually intractable; to overcome this, variational Bayes introduces an estimate of the posterior $q(z_{t})$. This is optimized to minimize the Kullback-Leibler divergence (KLD) between the estimated posterior $q(z_{t})$ and the true posterior $p(z_{t}|o_{t})$.
$D_{KL}[q(z_{t})||p(z_{t}|o_{t})]$ $=\int q(z_{t})\log\frac{q(z_{t})}{p(z_{t}|o_{t})}dz_{t}$
$=\int q(z_{t})\log\frac{q(z_{t})p(o_{t})}{p(z_{t},o_{t})}dz_{t}$
$=\int q(z_{t})\log\frac{q(z_{t})p(o_{t})}{p(z_{t})p(o_{t}|z_{t})}dz_{t}$ (S1)
$=F+\log{p(o_{t})}$ (S2)
The term $F$ here is the evidence free energy, equal to
$F_{t}=\underbrace{D_{KL}[q(z_{t})||p(z_{t})]}_{\text{Complexity}}-\underbrace{\mathbb{E}_{q(z_{t})}[\log p(o_{t+1}|z_{t})]}_{\text{Accuracy}}.$ (S3)
Since $p(o_{t})$ is constant for a given sensory observation, minimizing KLD is equivalent to minimizing $F_{t}$. Therefore, the optimal posterior estimate is:
$q^{*}(z_{t})=\underset{q(z_{t})}{\text{arg min }}F_{t}$ (S4)
In active inference, the agent minimizes expected free energy $G_{\tau}$ at a future time step $\tau\geq t+1$. This is the expected value of the evidence free energy under the predictive distribution of future outcomes (?).
$G_{\tau}$ $=\mathbb{E}_{p(o_{\tau}|z_{\tau})}[F]$
$=\mathbb{E}_{p(o_{\tau}|z_{\tau})}[\int q(z_{\tau})\log\frac{q(z_{\tau})}{p(o_{\tau},z_{\tau})}dz]$
$=\mathbb{E}_{p(o_{\tau}|z_{\tau})}[\mathbb{E}_{q(z_{\tau})}[\log\frac{q(z_{\tau})}{p(z_{\tau}|o_{\tau})}-\log p(o_{\tau})]].$ (S5)
Recalling that $q(z_{\tau}|o_{\tau})q(o_{\tau})=q(o_{\tau},z_{\tau})$, we approximate:
$G_{\tau}$ $\approx\mathbb{E}_{q(o_{\tau},z_{\tau})}[\log\frac{q(z_{\tau})}{q(z_{\tau}|o_{\tau})}-\log p(o_{\tau})]$
$=-\mathbb{E}_{q(o_{\tau},z_{\tau})}[\log\frac{q(z_{\tau}|o_{\tau})}{q(z_{\tau})}]-\mathbb{E}_{q(o_{\tau})}[\log p(o_{\tau})]$
$=-\underbrace{\mathbb{E}_{q(o_{\tau})}[\overbrace{D_{KL}[q(z_{\tau}|o_{\tau})||q(z_{\tau})]}^{\text{Bayesian Surprise}}]}_{\text{Epistemic Value or Mutual Information}}-\underbrace{\mathbb{E}_{q(o_{\tau})}[\log p(o_{\tau})]}_{\text{Extrinsic Value}}.$ (S6)
The first term, $I(z_{\tau},o_{\tau})=\mathbb{E}_{q(o_{\tau})}[D_{KL}[q(z_{\tau}|o_{\tau}]||q(z_{\tau})]]$, is the mutual information (or Bayesian surprise). This depicts expected information gain based on new sensory observation $o_{\tau}$, and can be expressed as:
$I(z_{\tau},o_{\tau})=\underbrace{H(z_{\tau})}_{\text{Shannon Entropy}}-\underbrace{H(z_{\tau}|o_{\tau})}_{\text{Conditional Entropy}}.$
The second term, $p(o_{\tau})$, represents log-likelihood of the preferred sensory observation. This is specified as the extrinsic reward designed by the experimenters. For the intrinsic value to reflect mutual information or information gain, and the extrinsic value to reflect expected free energy, is the same as the way shown by Friston’s group in the study of active inference (?, ?). Separating $o_{t}$ into $o_{t}$ and $a_{t}$, we rewrite the expected free energy as:
$G_{\tau}$ $=-\mathbb{E}_{q(o_{\tau},a_{\tau},z_{\tau})}[\log\frac{p(z_{\tau}|o_{\tau},a_{\tau})}{q(z_{\tau})}]-\mathbb{E}_{q(o_{\tau},a_{\tau})}[\log p(o_{\tau},a_{\tau})]$
$=-\mathbb{E}_{q(o_{\tau},a_{\tau},z_{\tau})}[\log\frac{p(z_{\tau},a_{\tau}|o_{\tau})}{q(z_{\tau})p(a_{\tau}|o_{\tau})}]-\mathbb{E}_{q(o_{\tau},a_{\tau})}[\log p(o_{\tau},a_{\tau})]$
$\approx-\mathbb{E}_{q(o_{\tau},a_{\tau},z_{\tau})}[\log\frac{q(z_{\tau}|o_{\tau})q(a_{\tau}|o_{\tau},z_{\tau})}{q(z_{\tau})p(a_{\tau}|o_{\tau})}]-\mathbb{E}_{q(o_{\tau},a_{\tau})}[\log p(o_{\tau},a_{\tau})]$
$=-\mathbb{E}_{q(a_{\tau}|o_{\tau},z_{\tau})q(o_{\tau})}[D_{KL}[q(z_{\tau}|o_{\tau})||q(z_{\tau})]]$
$-\mathbb{E}_{q(o_{\tau},z_{\tau})}[D_{KL}[q(a_{\tau}|o_{\tau},z_{\tau})||p(a_{\tau}|o_{\tau})]]$
$-\mathbb{E}_{q(o_{\tau},a_{\tau})}[\log p(o_{\tau},a_{\tau})].$ (S7)
Kawahara et al. developed a forward model $f_{w}(o_{\tau},a_{\tau})\rightarrow\widehat{o}_{\tau+1}$ which learns to predict the future sensory observation $o_{\tau+1}$ based $o_{\tau}$ and $a_{\tau}$ using a Bayesian Neural Network (BNN) (?). In this type of model, the network parameters $w_{\tau}$ are treated as random variables defined with gaussian distribution. These parameters serve as latent causes of observed sensory transitions and can be interpreted as random latent variables for the generative model. Therefore, $w_{\tau}$ corresponds to $z_{\tau}$.
Let the approximate posterior be defined as $q_{\psi}=\mathcal{N}(w_{\tau}|\mu,\sigma)$, with parameters $\psi=\{\mu,\sigma\}$. In this setting, the actor $\pi_{\phi}$ of a SAC can be trained to approximate $\pi_{\phi}(a_{\tau}|o_{\tau})\approx q(a_{\tau}|o_{\tau},w_{\tau})$. This allows rewriting the expected free energy as:
$G(o_{\tau},a_{\tau})$ $=-\mathbb{E}_{q(a_{\tau}|o_{\tau},w_{\tau})q(o_{\tau})}[D_{KL}[q(w_{\tau}|o_{\tau})||q(w_{\tau})]]$
$-\mathbb{E}_{q(o_{\tau},a_{\tau})}[D_{KL}[\pi_{\phi}(a_{\tau}|o_{\tau})||p(a_{\tau}|o_{\tau})]]$
$-\mathbb{E}_{q(o_{\tau},a_{\tau})}[\log p(o_{\tau},a_{\tau})].$ (S8)
Let us interpret the prior preference $\log p(o_{\tau},a_{\tau})$ as the extrinsic reward $r(s_{\tau},a_{\tau})$, where $s_{\tau}$ is the true environmental state. Bring focus to the current time step by setting $\tau=t$. Because the forward model trains to predict $o_{t+1}$, we can further rewrite the expected free energy as:
$G(o_{t},a_{t})$ $=-D_{KL}[q_{\psi}(w_{t}|o_{t+1})||q_{\psi}(w_{t})]-\log p(o_{t},a_{t})$
$-D_{KL}[\pi_{\phi}(a_{t}|o_{t})||p(a_{t}|o_{t})]$
$=-D_{KL}[q_{\psi}(w_{t}|o_{t+1})||q_{\psi}(w_{t})]-\log p(o_{t},a_{t})$
$-\int\pi_{\phi}(a_{t}|o_{t})\log\pi_{\phi}(a_{t}|o_{t})da_{t}+\int\pi_{\phi}(a_{t}|o_{t})\log p(a_{t}|o_{t})da_{t}$
$=-\underbrace{D_{KL}[q_{\psi}(w_{t}|o_{t+1})||q_{\psi}(w_{t})]}_{\text{Curiosity}}-\underbrace{r(s_{t},a_{t})}_{\text{Extrinsic Reward}}-\underbrace{\mathcal{H}(\pi_{\phi}(a_{t}|o_{t}))}_{\text{Entropy}}-\underbrace{\mathbb{E}_{\pi_{\phi}(a_{t}|o_{t})}[\log p(a_{t}^{\ast}|o_{t})]}_{\text{Imitation}}$ (S9)
Because $w_{\tau}$ represents the robot’s probabilistic knowledge of their environment, the first term of Eq. Free energy principle, Active Inference, and Kawahara Model can be said to represent the robot’s gain in knowledge based on information acquired in a new sensory observation.
In summary, the forward model is trained to minimize the evidence free energy $F$ (Eq. S3) by accurately reconstructing sensory observations and minimizing posterior complexity based on past experiences. Meanwhile, the actor-critic pair is trained to minimize expected free energy $G$, which includes an inverted complexity term (i.e., curiosity) and entropy to encourage exploration. This leads to emergent tension in an adversarial relationship: the actor is encouraged to maximize information gain by increasing the KL divergence between prior and posterior, which the forward model trains to minimize that same term. This establishes a dynamic push-pull effect, driving self-organized exploration. Please note that the imitation term in Eq. Free energy principle, Active Inference, and Kawahara Model depends on external demonstrations or expert policies; this term is ignored in our study, which focuses on self-exploration.
From this formulation of expected free energy, the $Q$-value can be updated as:
$Q(t)$ $=r_{t}+\eta D_{KL}[q_{\psi}(w_{t}|o_{t+1})||q_{\psi}(w_{t})]+$
$\gamma(1-done_{t})\mathbb{E}_{o_{t+1}\sim D,a_{t+1}\sim\pi_{\phi}}[Q_{\bar{\theta}}(o_{t+1},a_{t+1})]+\alpha\mathcal{H}(\pi_{\phi}(a_{t+1}|o_{t+1}))$ (S10)
Here, $\eta>0$ and $\alpha>0$ are hyperparameter weighting the intrinsic reward based on the curiosity and the motor entropy, respectively.
In our experiments, each episode ended after 30 steps, or terminated earlier if the agent successfully executed the command. Completed episodes are stored in a recurrent replay buffer, which can hold up to 256 episodes. When the buffer is full, the buffer discards the oldest episodes to accommodate new episodes. To ensure uniform episode length, all episodes were padded to 30 steps with empty transitions. Hence, transitions are stored with the form $\{o_{t},a_{t},r_{t},o_{t+1},done_{t},mask_{t}\}$, where $mask_{t}=1$ for real transitions, and $mask_{t}=0$ for empty transitions added for padding. After each episode, a batch of 32 episodes was sampled from the buffer and used to train the forward model, actor, and critics. During training, loss terms were multiplied by $mask_{t}$, removing the influence of empty transitions.
Implementation details
Vision
The robot visually senses the environment in the direction the robot faces with a $16x16x4$ image, with the four channels being red, green, blue, and distance. See Figure S1.
Figure S1: The agent’s vision, $\bm{o_{t,v}.}$ The robot is facing a magenta cone and a green pillar. The robot also sees part of its hand. The image on the left depicts the red, green, and blue channels. The image on the right depicts the distance.
In our proposed model, in order to make the estimated posterior for visual sensations, images are flattened and encoded using a linear neural network with Parametric Rectified Linear Unit activation (PReLU). To generate a prediction of the next image, $h^{q}_{t}$ and $a^{enc}_{t}$ are concatenated and decoded with another linear neural network, shaped into a $16x16x4$ tensor, and finished with a convolutional layer. See details in table S2.
Layer Type Activation Details
Encoder, $enc_{v}$
1 Flatten Shape (16, 16, 4) to shape (1024).
2 Linear PReLU To shape (128).
Decoder, $dec_{v}$
1 Linear BatchNorm2d, PReLU From shape (264) to shape (8 * 8 * 64).
2 Reshaping To shape (8, 8, 64).
3 CNN Tanh Kernel size 3, reflective padding 1. To shape (8, 8, 8).
4 Pixel Shuffle To shape (16, 4, 4).
Table S2: Encoder and decoder of agent’s visual sensations, $\bm{o_{t,v}}$.
Touch
The second part of the sensory observation is the tactile sensation of touch. This is represented by a one value between 0 and 1 for each of the 16 sensors. Each value is equal to the fraction of time in the previous step during which the respective sensor was in contact with an object. See Figure S2.
Figure S2: The agent’s sensors for tactile sensations of touch, $\bm{o_{t,ta}}$. The robot has 16 sensors, which are planes on the surface of the robot’s body, arm, and hand. The camera and wheels are marked just for clarity.
In our proposed model, in order to make the estimated posterior for tactile sensation, the tensor is encoded using a linear neural network with PReLU. To generate a prediction of the next tactile sensation, $h^{q}_{t}$ and $a^{enc}_{t}$ are concatenated and decoded with another linear neural network. See details in table S3.
Layer Type Activation Details
Encoder, $enc_{ta}$
1 Linear BatchNorm2d, PReLU From shape (16) to shape (20).
Decoder, $dec_{ta}$
1 Linear BatchNorm2d, TanH From shape (264) to shape (16). Result added to 1 and divided by 2 for values between 0 and 1.
Table S3: Encoder and decoder of agent’s tactile sensations, $\bm{o_{t,ta}}$.
Proprioception
The third part of the sensation is the angle and velocity of the arm’s joints. This consists of a tensor with four values between 0 and 1: two joint angles and two joint velocities. Each value is the normalized proportion of the respective variable between its minimum and maximum range.
In our proposed model, in order to make the estimated posterior for sensation of proprioception, the tensor is encoded using a linear neural network with PReLU. To generate a prediction of the next proprioception, $h^{q}_{t}$ and $a^{enc}_{t}$ are concatenated and decoded with another linear neural network. See details in table S4.
Layer Type Activation Details
Encoder, $enc_{po}$
1 Linear BatchNorm2d, PReLU From shape (4) to shape (4).
Decoder, $dec_{po}$
1 Linear BatchNorm2d, TanH From shape (264) to shape (4). Result added to 1 and divided by 2 for values between 0 and 1.
Table S4: Encoder and decoder of agent’s sensation of proprioception, $\bm{o_{t,p}}$.
Voices
The fourth and fifth parts of the sensation are the command voice and the tutor-feedback voices, which were described briefly in the Results section. Both voices are sequences of one-hot vectors. Table S5 displays the 18 terms (including silence) and their indexes in the one-hot vectors. For example, the command “Watch the Red Pillar” is represented by
$[0,0,0,0,0,0,0,1,0,0,0,0,0,0,0,0,0,0]$
$[0,0,0,0,0,0,0,0,0,0,0,0,0,1,0,0,0,0].$
If the robot has not performed any action, then the feedback voice is only one one-hot vector indicating silence:
. (S12)
English Word Indexes
Index Word
0 (Silence)
1 Watch
2 Be Near
3 Touch the Top
4 Push Forward
5 Push Left
6 Push Right
Index Word
7 Red
8 Green
9 Blue
10 Cyan
11 Magenta
12 Yellow
13 Pillar
14 Pole
15 Dumbbell
16 Cone
17 Hourglass
Table S5: English words and their indexes in one-hot vectors.
The robot’s forward model’s encoding of these two voices has two parts. The first part of the encoding is an embedding and recurrent neural network. This part is identical for the command voice and the feedback voices, ensuring that tokens are interpreted consistently across sources. Note that this RNN is “nested” within the forward model’s RNN, such that each of the robot’s steps includes three steps of interpreting voices. See figure S3. In the second part of the encoding, outputs from the first part of the encoding are processed with unique linear layers to produce separate estimated posteriors. To generate a prediction of the next voices, $h^{q}_{t}$ and $a^{enc}_{t}$ are concatenated and decoded using two separate recurrent neural networks for the command voice and feedback voice. See details in table S6.
Layer Type Activation Details
Encoder part one, $enc_{w}$ (shared by command voice and feedback voice)
1 Embedding PReLU From shape (Sequence-length, 18) to shape (Sequence-length, 8).
2 Linear PReLU To shape (Sequence-length, 64).
3 GRU PReLU To shape (64).
4 Linear PReLU To shape (256).
Decoders, $dec_{cw}$ and $dec_{fw}$
1 Linear BatchNorm2d, PReLU From shape (264) to shape (192).
2 Reshaping To shape (3, 64).
3 GRU PReLU To shape (3, 64).
4 Linear To shape (3, 17).
Table S6: Encoder and decoder of agent’s voice sensation, $\bm{o_{t,cw}}$ and $\bm{o_{t,fw}}$.
Figure S3: Recurrent step shared by command voice and feedback voice.
Motor Command Encoder
For usage in the forward model, the robot’s motor commands $a_{t}$ are encoded into $a^{enc}_{t}$ with a linear neural network with PReLU. See details in table S6.
Layer Type Activation Details
Encoder, $enc_{a}$
1 Linear PReLU From shape (4) to shape (8).
Table S7: Encoding motor command for forward model.
Constraints in Performing Actions
In each step, the robot can only perform one of the six actions. This is implemented using definitions of actions and action prioritization. The actions Watch, Be Near, and Touch the Top cannot be performed simultaneously because of requirements regarding distance from the object and touching the object. The actions Push Left and Push Right cannot be performed simultaneously because of the directions of movements. If the robot satisfies the requirements for Touch the Top, we reject the actions Push Forward, Push Left, or Push Right. If the robot is performing Push Forward and Push Left or Push Right, we allow only the actions with the greatest distance pushed.
Details of Experiment Design
10 robots trained in each way described in the Results section: no curiosity, sensory-motor curiosity, and all curiosity. The robots trained for 60000 epochs. In each epoch, the robot performed one episode which was saved in its recurrent replay buffer. Then the robot trained with a batch of 32 of its saved episodes.
Experiment 1
Experiment 1 tests the effects of curiosity. We trained robots with three levels of curiosity: no curiosity, sensory-motor curiosity, and all curiosity. In table S8, we share the value of the $\eta$ hyperparameters for each of the four parts of the sensory observation which may be explored. These represent the relative importance of each part of the sensory observation in the robot’s curiosities.
Name $\eta_{vision}$ $\eta_{touch}$ $\eta_{proprioseption}$ $\eta_{feedback}$
No Curiosity 0 0 0 0
Sensory-Motor Curiosity .05 2 .1 0
Complete Curiosity .05 2 .1 .3
Table S8: Hyperparameters for three types of agents.
We measured the success-rates of these three types of robots in the six types of actions. The plots in figure 3 show the rolling average of success-rates of the three types robots from the beginning of training to the end of training after 60000 epochs, with 99% confidence intervals. Specifically, the plots show results of the robots regarding the goals with combinations of action, color, and shape which the robots were not shown in training, testing for the ability to generalize vocabulary and syntax to unlearned combinations.
As we predicted in hypothesis $i$, robots with no curiosity performed the worst, with approximately 25% success-rate; robots with curiosity for sensory-motor observations performed better, with approximately 75% success-rate; and robots with curiosity for sensory-motor observations and the feedback voice are the best, with approximately 90% success-rate. As we predicted in hypothesis $ii$, the robot’s ability to perform simpler actions develop earliest, and the robot’s ability to perform more complex actions develop later, having required the simpler actions as prerequisites. Merely watching the object appears to be the simplest, developing earliest, while pushing object the object to the left or right appear to be the most complex, developing later.
Experiment 2
Experiment 2 tests the relationship between success-rates with learned goals and unlearned goals, specifically by robots using all curiosity. See figure 4. The left column shows success-rate plots of robots with learned actions, while the right column shows success-rate plots of robots with unlearned actions. The first row shows results for robots using the complete vocabulary: 6 actions, 6 colors, and 5 shapes. The second and third row show results for robots trained with smaller vocabularies. In each of the three situations, the robots are trained with one third of the possible goals, and tested with the other two thirds.
As we predicted in hypothesis iiiiii, the robot’s success-rates with learned actions initiates earlier than its success-rates with unlearned actions. This suggests pairing sentences of words precedes generalization with compositionality. As we predicted in hypothesis $iv$, larger vocabularies lead to faster generalization. All three collections of robots had success-rates of approximately 100% with learned actions. Robots which were trained with 60 of the 180 possible goals with 6 actions, 6 colors, and 5 shapes had success-rates of approximately 90% with unlearned tasks. Robot which were trained with 25 of the 75 possible goals with 5 actions, 5 colors, and 3 shapes had success-rates of approximately 50% with unlearned tasks. And robots which were trained with 16 of the 48 possible goals with 4 actions, 4 colors, and 3 shapes has success-rates of approximately 30% with unlearned tasks. The ability to generalize quickly is enhanced with the size of the vocabulary in use.